OpenStudy (anonymous):

Solve x2 - 4x = 12 by completing the square. Please show work.

OpenStudy (anonymous):

I have to add 4 to each side i know that so then i get x^2-4x+4=16 then after that i think i need to factor the left side and i get (x-2)^2=16

OpenStudy (anonymous):

then i think i need to square root both sides and i get x-2=4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so after this i should add 2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but how do i get 2 answers?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so x=6

OpenStudy (anonymous):

x-2=\[\pm4 \]

OpenStudy (anonymous):

x-2=4 or x-2=-4

OpenStudy (anonymous):

x=6 or x=-2

OpenStudy (anonymous):

oh ok so x-2=4 or x-2=-4 then add two to both sides to get x=6 or x=-2?

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me on my question lol

OpenStudy (anonymous):

i can try